OpenStudy (anonymous):

help . solve 2^3x+5=128

OpenStudy (ajprincess):

128=2^7 3x+5=7 solve for x. Does that help? @msnov93

OpenStudy (anonymous):

kinda , still confused

OpenStudy (ajprincess):

3x+5=7 u have to isolate x to solve for x. 1.Subtract 5 from both sides. 2.Divide both sides by 3

OpenStudy (anonymous):

thats it ?

OpenStudy (ajprincess):

yup

OpenStudy (anonymous):

k